can you help me solve this. linear system equations by elimination
x + y + z =4
x - y + 2z=1
x - y - 3z = -4

Answer by Alan3354(69443)

x + y + z =4
x - y + 2z=1
x - y - 3z = -4
============
All the x coeff's are 1, eliminate x first by subtracting 1 of the eqns from the other 2.
----------
If you subtract eqn 3 from 2, both x & y are eliminated.